OpenStudy (anonymous):

How do I put y=-5x^2-40x-6 in intercept form

OpenStudy (daniellelovee):

can it be divided into 2 equations? Because that is the only possible way that this equation is making sense to me

OpenStudy (anonymous):

I think so

OpenStudy (daniellelovee):

yeah this question seems easy but is actually not

OpenStudy (misty1212):

Hi! what is "intercept form"?

OpenStudy (daniellelovee):

y=mx+b she/he meant slope intercept form

OpenStudy (misty1212):

lol no

OpenStudy (misty1212):

it is not a line, so it will not look like \(y=mx+b\)

OpenStudy (daniellelovee):

lol yeah well i came here to answer because it appeared easy and now I want to know how you solve so yeah... if you know how then explain please :)

OpenStudy (misty1212):

maybe it means "factor"

OpenStudy (misty1212):

but this one does not factor with integers, so who knows?

OpenStudy (daniellelovee):

yeah I tried to divide it into (x-a)(x-b)=c but it didnt gave an answer either

OpenStudy (anonymous):

Slope intercept form only applies to straight lines.\[y=-5x^2-40x-6\]is not a straight line, and so it can't be represented using slope intercept form.

OpenStudy (anonymous):

You can factor\[y=-5x^2-40x-6\]to become:\[y=-(5x^2+40x+6)=-(5x+40x+6)=-[(5x+0)(x+8)+6]=-(5x+0)(x+8)-6\]

OpenStudy (daniellelovee):

@mathmate

OpenStudy (mathmate):

For a quadratic, intercept form means f(x)=a(x-p)(x-p) where p, q are the roots of the function.  As @misty1212 pointed out, this one does not have rational factors, so it is less obvious.

OpenStudy (mathmate):

To work out this problem, we need to solve the quadratic and find the two irrational roots p,q. Using the quadratic formula for f(x)=\(-5x^2-40x-6\) a=-5, b=-40 c=-6 x=\(\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\) =\(\dfrac{40\pm\sqrt{40^2-4(-5)(-6)}}{2(-5)}\) =\(\dfrac{20\pm\sqrt{370}}{5}\) Putting the roots into the intercept form gives f(x)=\(-\dfrac{1}{5}(5x+(\sqrt{370}+20))(5x-(\sqrt{370}-20))\) which is the final intercept form. You should always expand the form to confirm that it is indeed f(x)=\(-5x^2-40x-6\) I will leave that to you as part of the exercise. I have exceptionally provided the final answer, partly because the arithmetic is tedious, and partly because I suspect there is a typo in the question. So whenever we come to a result like this, it is always a good idea to double check the question, in exams as well as in exercises.